Great Falls Company makes two products, Wool Gloves and Wool Mittens. They are initially processed from the same raw material and then, after split-off, further processed separately. Additional information is as follows:

Gloves Mittens Total
Final Sales Price $ 13,200 $ 10,620 $ 23,820
Joint Costs Prior to Split-Off Point ? ? $ 9,012
Costs Beyond Split-Off Point $ 4,400 $ 4,400 $ 8,800


What are the joint costs allocated to Gloves and Mittens assuming Great Falls uses the estimated net realizable value approach?

Gloves Mittens
A. $ 4,506 $ 4,506
B. $ 4,840 $ 4,172
C. $ 5,280 $ 3,732
D. $ 5,440 $ 3,572

Multiple Choice

  • Option C

  • Option B

  • Option D

  • Option A
Answer #1
Gloves Mittens Total
NRV (13,200-4,400) = 8,800 (10,620-4,400) = 6220 15,020
Percent of total

= (8,800/15,020)

= 0.58588

= (6220/15,050)

= 0.41411

1

Joint cost allocated to

Gloves = 9012 * 0.58588 = 5280

Mittens = 9012 * 0.41411 = 3732

Option C is the answer (don't get confused. Select the first one)